Qd= 30 - 5P

Qs= -10 + 3P


1
Expert's answer
2022-03-22T11:46:02-0400

Qd=Qs

30-5P=-10+3P

40=8P

P=5

Qd= 30-5(5)

Qd=5

Qs=-10+3(5)

Qs=5